The magnifying power of a magnifying glass of power 12 dioptre is ?

Grade:Upto college level

2 Answers

Sameee Sharma
5 years ago
Given, p= +12 diopter
          So, f= 1/p
                  = 1/12 m = 0.083m
Here, D= 25cm = 0.25m
Now by , M= 1+ D/f
                    = 1+ 0.25/0.083
                    = 1+ 3.01
                    = 4.01~ 4
Rohan kumar
2 years ago
Given, p= +12 diopter
                  = 1/12 m = 0.083m
 
             Here, D= 25cm = 0.25m
 
Now by , M= 1+ D/f
                    = 1+ 0.25/0.083
                    = 1+ 3.01
                    = 4.01.
            So, = 4  answer.
 
Thanks .